### v1

@@ Abstract
Keywords:
Psychiatric Diagnosis Brain Graph Representation Self-supervised Learning Structure Semantics fMRI Graph Learning.
@@ 3 The Proposed Method
The overall framework of SAM-BG is shown in Figure 1. It consists of two main phases: (i) Structural semantic extraction, where an edge masker is trained to identify substructures containing task-relevant structural semantics; (ii) SSL with structure-sematic preservation, where augmentations are applied only to non-salient edges while retaining the extracted substructures, thus enhancing representation learning by maintaining critical structural semantics.
@@ 3.1 Structural Semantic Extraction
We represent a brain network as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{A},\mathbf{X})$, where $\mathcal{V}=\{v_{1},v_{2},\dots,v_{N}\}$ denotes $N$ ROIs, $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the adjacency matrix computed via Pearson correlation [21], and $\mathbf{X}\in\mathbb{R}^{N\times D}$ is the node feature matrix with each row $\mathbf{x}_{i}$ representing the feature vector of ROI $v_{i}$. An edge $e_{i,j}$ exists if $a_{i,j}$ indicates significant functional connectivity. Since functional connections are critical for diagnosing psychiatric disorders [7], we introduce an edge masker to extract discriminative substructures from $\mathbf{A}$ as structural prior knowledge.
Specifically, the masker consists of a multi-layer perceptron (MLP) with sigmoid activation, which encodes the upper triangular part of $\mathbf{A}$ to generate a probabilistic edge-selection matrix $\mathbf{P}\in[0,1]^{N\times N}$, mirrored to preserve symmetry. To enable discrete but differentiable edge sampling, Gumbel-Softmax reparameterization [15] is applied to binarize $\mathbf{P}$ into $\mathbf{P}_{\text{edge}}\in\{0,1\}^{N\times N}$. The final subgraph is then obtained by masking the original adjacency as $\mathbf{A}_{\text{sub}}=\mathbf{A}\odot\mathbf{P}_{\text{edge}}$. Then, the masker is trained under the Information Bottleneck (IB) principle [28], as illustrated in Figure 1 (left), aiming to retain label-relevant semantics while discarding redundant topology. Specifically, the objective is to minimize the mutual information (MI) between $\mathcal{G}_{\text{sub}}$ and $\mathcal{G}$ to encourage compression, while maximizing the MI between $\mathcal{G}_{\text{sub}}$ and the label $y$ to ensure discriminability. The objective is formally expressed as:
@@ 3.2 Self-supervised Learning with Structure-sematic Preservation
As shown in Figure 1 (right), we propose a representation learning module consisting of a structure-aware graph augmentation strategy with a tailored learning objective to learn semantically meaningful representations for data-efficient diagnosis of psychiatric disorders.
@@
An effective data augmentation strategy should increase diversity while preserving semantic consistency. To achieve this, we leverage a trained edge masker to transfer structural priors from limited labeled data to large-scale unlabeled data, thereby generating biologically plausible augmented brain graphs that support effective representation learning. Specifically, the masker identifies salient substructures $\mathcal{G}_{\text{sub}}$ from an unlabeled graph $\mathcal{G}$. Augmentation is then performed by introducing controlled perturbations into the complementary non-salient region $\Delta\mathcal{G}=\mathcal{G}\backslash\mathcal{G}_{\text{sub}}$, where edges are randomly removed under a uniform distribution. This process yields two structure-aware augmented views, $\mathcal{G}_{a}$ and $\mathcal{G}_{b}$, which preserve essential semantic structures while ensuring sufficient diversity.
@@
To capture shared structural semantics across views, SAM-BG employs canonical correlation analysis (CCA) [3] as the core learning objective. Specifically, the two augmented graphs, $\mathcal{G}_{a}$ and $\mathcal{G}_{b}$, are passed through a shared GNN encoder, which is initialized from the pre-training phase. The encoder outputs view-specific embeddings $\mathbf{Z}_{a}$ and $\mathbf{Z}_{b}$, which are aligned in the latent space by maximizing their correlation via a CCA-based loss. The learning objective is defined as:
@@
where the $\lambda$ is the trade-off coefficient, and $\mathbf{I}$ is an identity matrix. The first term of the objective aligns the embeddings of the two augmented graphs in a shared latent space, promoting perturbation-invariant representations. The second term enforces decorrelation across embedding dimensions, preserving view-specific features and preventing representation collapse.
During inference, unlabeled brain graphs $\mathcal{G}$ are first processed by the pre-trained edge masker to retain key substructures. These structure-aware augmented graphs are then passed through the trained GNN encoder to generate consistent representations for downstream psychiatric disorder classification, ensuring interpretable and clinically relevant features from unlabeled data.
@@ 4.1 Experiment Setup
We conducted experiments on two publicly available medical imaging datasets: ABIDE  [^1] https://fcon_1000.projects.nitrc.org/indi/abide/ and ADHD-200  [^2] https://fcon_1000.projects.nitrc.org/indi/adhd200/. Both contain resting-state fMRI data commonly used in psychiatric research. For ABIDE, we used 743 subjects, 364 with autism spectrum disorder (ASD) and 379 healthy controls (HC), aged 6–64, with balanced age and gender distribution. For ADHD-200, we selected 459 participants, including 230 with ADHD and 229 typically developing (TD) individuals, aged 7–21. Brain networks were constructed using the Automated Anatomical Labeling (AAL) atlas, which defines 116 regions of interest ROIs as nodes. Pearson correlation was used to compute edges between ROI time series, resulting in weighted brain graphs.
@@
We compare SAM-BG with several representative baselines, including both supervised and SSL methods. For supervised learning, we use BrainGNN [11], which applies ROI-aware pooling, BrainGB [4], and a standardized benchmark. For SSL, we include GraphCL [27], GCA [32], MA-GCL [8], and CCA-SSG [30], which differ in their augmentation and view generation strategies. We adopt the official settings for BrainGNN and BrainGB. All SSL baselines use their original implementations with edge perturbation and 20% labeled data for fine-tuning, simulating low-supervision scenarios.
@@
Our model is implemented using PyTorch and trained on an NVIDIA Tesla P100 GPU. During pre-training, we leverage 20% of the labeled data to guide masker training. For downstream tasks, the dataset is split into 80% for fine-tuning, 10% for validation, and 10% for testing, with the test set strictly excluded from any supervised training during pre-training. The hyperparameters are configured as follows: the MI trade-off coefficient $\beta$ is set to 0.01 for ABIDE and 0.001 for ADHD, while the weight of the decorrelation loss term $\lambda$ is set to $1\text{e}^{-4}$ for ABIDE and 0.001 for ADHD. We conduct 5-fold cross-validation and repeat experiments five times with different random seeds. Results are reported as mean and standard deviation of accuracy (ACC), the area under the ROC curve (AUC), recall, and F1-score.
@@ 4.2 Experimental Results and Discussion
Table 1 shows the comparison between our method and baselines on the ABIDE and ADHD datasets. Overall, supervised methods generally outperform most SSL baselines, highlighting the value of label information in graph classification. BrainGB achieves the highest F1-score among supervised methods, showing strong performance under full supervision. Among SSL methods, CCA-SSG performs well due to its non-contrastive design, which better preserves structural-semantic features. Our method, SAM-BG, consistently outperforms all baselines on both datasets. On ABIDE, it improves ACC by 4.1%, AUC by 3.4%, and F1-score by 3.7% compared to the best baseline, although its recall is slightly lower than A-GCL. On ADHD, it achieves 3.3% higher ACC, 0.3% higher recall, and 3.7% higher F1-score, with a marginally lower AUC than A-GCL. Notably, SAM-BG demonstrates more substantial improvements on the ABIDE dataset. This may indicate that disease-related connectivity patterns are more discriminative in the brain networks of individuals with autism, and that SAM-BG’s two-stage training strategy effectively captures these structural and semantic features to deliver superior results.
@@
The Effectiveness of Structure-aware SSL in Data-efficient Settings.
@@ 5 Conclusion
We propose SAM-BG, a novel framework that leverages structural priors to enable biologically meaningful brain graph representation learning under limited supervision. SAM-BG first trains an edge masker on a small labeled subset to extract salient connections, capturing key structural semantics. It then generates structure-aware augmentations by perturbing only non-essential edges, preserving semantic integrity for robust SSL. Experiments show that SAM-BG significantly improves psychiatric disorder detection in low-data settings and offers interpretable insights into brain connectivity, highlighting its potential for real-world neuroscience applications.

### v2

@@ 3 The Proposed Method
The overall framework of SAM-BG is shown in Figure 1. It consists of two main phases: (i) Structural semantic extraction, where an edge masker is trained to identify substructures containing task-relevant structural semantics; (ii) SSL with structure-sematic preservation, where augmentations are applied only to non-salient edges while retaining the extracted substructures, thus enhancing brain network representations by maintaining critical structural semantics.
@@
3.1 Problem Definition
We represent a brain network as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{A},\mathbf{X})$, where $\mathcal{V}=\{v_{1},v_{2},\dots,\\ v_{N}\}$ denotes the set of $N$ ROIs, $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the adjacency matrix computed via Pearson correlation [22], and $\mathbf{X}\in\mathbb{R}^{N\times D}$ is the node feature matrix with each row $\mathbf{x}_{i}\in\mathbb{R}^{D}$ representing the feature vector of ROI $v_{i}$. An edge $e_{i,j}\in\mathcal{E}$ exists if $a_{i,j}\in\mathbf{A}$ indicates significant functional connectivity. Given a small labeled dataset $\{\mathcal{G}^{i},y^{i}\}_{i=1}^{S}$ and a large unlabeled set $\{\mathcal{G}^{j}\}_{j=1}^{U}$, where $U\gg S$ and $y\in\{0,1\}$, our goal is to learn biologically meaningful brain graph representations $\mathbf{Z}\in\mathbb{R}^{N\times D}$, which are further used for downstream tasks.
@@ 3.2 Structural Semantic Extraction
Since brain connections (edges) are pivotal for diagnosing and interpreting the neural mechanisms of psychiatric disorders [10, 7], we introduce an edge masker to distill discriminative substructures as structural priors (Figure 2). For a given labeled brain network $\mathcal{G}$, the edge masker operates directly on $\mathbf{A}$, encoding it via a two-layer multi-layer perceptron (MLP) network followed by a sigmoid activation function to generate edge selection probabilities $\mathbf{P}\in\mathbb{R}^{N\times N}$. Each element $p_{i,j}\in[0,1]$ denotes the selection probability of edge $e_{i,j}$. To impose sparsity while avoiding over-pruning, we divide the probability matrix $\mathbf{P}$ into $K$ submatrices, each containing ${N\times N}/{K}$ edges. Within each group, the Gumbel–Softmax reparameterization [16] selects one edge among $K$ consecutive candidates, yielding a differentiable, structurally coherent mask. The grouped masks are reshaped back to $N\times N$ size and binarized to form the final edge selection mask $\mathbf{P}_{\text{edge}}$. The substructure of brain network $\mathcal{G}_{\text{sub}}$ is obtained by element-wise masking of the original graph: $\mathbf{A}_{\text{sub}}=\mathbf{A}\odot\mathbf{P}_{\text{edge}}$.
Then, the masker is trained under the Information Bottleneck (IB) principle [29], as illustrated in Figure 1 (left), aiming to retain label-relevant semantics while discarding redundant topology. Specifically, the objective is to minimize the mutual information (MI) between $\mathcal{G}_{\text{sub}}$ and $\mathcal{G}$ to encourage compression, while maximizing the MI between $\mathcal{G}_{\text{sub}}$ and the label $y$ to ensure discriminability. The objective is formally expressed as:
@@
Figure 2: Framework of the subgraph generation process. The input graph $\mathcal{G}$ is processed by the masker to sample a subset of edges and construct a subgraph $\mathcal{G}_{\text{sub}}$ that preserves key structural information.
@@ 3.3 Self-supervised Learning with Structure-sematic Preservation
As shown in Figure 1 (right), we propose an SSL-based representation learning module, which incorporates structural priors to mitigate the information loss commonly introduced by random perturbations in existing methods. The framework comprises a structure-aware graph augmentation strategy with a tailored learning objective to learn semantically meaningful representations for data-efficient diagnosis of psychiatric disorders.
@@
[…]
@@
After generating structurally meaningful augmented brain graphs, we adopt an SSL framework to learn graph-level representations. To capture shared structural semantics across views, SAM-BG employs canonical correlation analysis (CCA) [3] as its core learning objective. CCA encourages the model to learn perturbation-invariant yet semantically aligned embeddings by maximizing the correlation between the augmented view representations. Specifically, the two augmented brain graphs, $\mathcal{G}_{a}$ and $\mathcal{G}_{b}$, are fed into a shared GNN encoder, which is initialized using the encoder pre-trained during the structure masker phase. This encoder generates view-specific embeddings $\mathbf{Z}_{a}$ and $\mathbf{Z}_{b}$, which are then optimized via a CCA-based loss to align the representations in the latent space. The learning objective is defined as:
@@
where the $\lambda$ is the trade-off coefficient, and $\mathbf{I}$ is an identity matrix. The first term of the objective aligns the embeddings of the two augmented graphs in a shared latent space, promoting perturbation-invariant representations. The second term enforces decorrelation across embedding dimensions, preserving view-specific features and preventing representation collapse. By jointly optimizing these two terms, the trained GNN encoder can effectively learn meaningful and discriminative representations $\mathbf{Z}$ from any given brain network $\mathcal{G}$. These representations can then be applied to various downstream tasks, such as psychiatric disorder classification or biomarker discovery. The overall training procedure of SAM-BG is summarized in Algorithm 1.
Algorithm 1 Training Procedure of the SAM-BG Model
1:Input: Labeled brain network dataset $\{\mathcal{G}^{i},y^{i}\}_{i=1}^{S}$, Unlabeled brain network dataset $\{\mathcal{G}^{j}\}_{j=1}^{U}$, Training epochs $T$
2:Output: Optimized model parameters $\hat{\theta}_{\text{GNN}},\hat{\theta}_{\text{Masker}}$
3:Initialize: Model parameters $\theta_{\text{GNN}},\theta_{\text{Masker}}$
4:Phase I: Structural Semantic Extraction Phase (supervised)
5:for epoch =1=1 to $T$ do
6:  for each labeled sample $i=1$ to $S$ do
7:   Compute edge mask probability $\mathbf{P}^{i}=f_{\theta_{\text{Masker}}}(\mathcal{G}^{i})$
8:   Binarize $\mathbf{P}^{i}$ to obtain substructure $\mathcal{G}_{\text{sub}}^{i}$
9:   Obtain representations $\mathbf{Z}^{i}=f_{\theta_{\text{GNN}}}(\mathcal{G}^{i})$ and $\mathbf{Z}_{\text{sub}}^{i}=f_{\theta_{\text{GNN}}}(\mathcal{G}_{\text{sub}}^{i})$
10:   Compute loss $\mathcal{L}_{\text{pre}}$ (Equation 2)
11:   Update $\theta_{\text{GNN}}$, $\theta_{\text{Masker}}$ via gradient descent
12:  end for
13:end for
14:Return: Trained $\hat{\theta}_{\text{Masker}}$ and warm-start ${\theta}_{\text{GNN}}$
15:Phase II: Self-supervised Representation Learning Phase
16:for epoch =1=1 to $T$ do
17:  for each unlabeled sample $j=1$ to $U$ do
18:   Extract substructure $\mathcal{G}_{\text{sub}}^{j}$ using trained masker $f_{\hat{\theta}_{\text{Masker}}}(\mathcal{G}^{j})$
19:   Construct augmented graphs $\mathcal{G}_{a}^{j}$ and $\mathcal{G}_{b}^{j}$ by perturbing $\Delta\mathcal{G}^{j}=\mathcal{G}^{j}\backslash\mathcal{G}_{\text{sub}}^{j}$
20:   Obtain representations $\mathbf{Z}_{a}^{j}=f_{\theta_{\text{GNN}}}(\mathcal{G}_{a}^{j})$ and $\mathbf{Z}_{b}^{j}=f_{\theta_{\text{GNN}}}(\mathcal{G}_{b}^{j})$
21:   Compute loss $\mathcal{L}_{\text{CCA}}$ (Equation 3)
22:   Update $\theta_{\text{GNN}}$ via gradient descent
23:  end for
24:end for
25:Return: $\hat{\theta}_{\text{GNN}},\hat{\theta}_{\text{Masker}}$
@@ 4.1 Experiment Setup
We conducted experiments using two real-world medical datasets: Autism Brain Imaging Data Exchange (ABIDE)  [^1] https://fcon_1000.projects.nitrc.org/indi/abide/ and Attention Deficit Hyperactivity Disorder (ADHD-200)  [^2] https://fcon_1000.projects.nitrc.org/indi/adhd200/ datasets. Both contain resting-state fMRI data commonly used in psychiatric research. For ABIDE, we selected a subset of 743 participants, comprising 364 ASD patients (ages 7–24) and 379 HC individuals (ages 6–22), with balanced age and gender distribution. For ADHD-200, we utilized a sample of 459 participants from this dataset, consisting of 229 typically developing (TD) individuals and 230 children and adolescents diagnosed with ADHD (ages 7–21).
Data Preprocessing
We preprocess resting-state fMRI into graph-ready brain networks. The pipeline includes slice-timing correction, head-motion correction, spatial normalization, and Gaussian smoothing, implemented with the Graph Theoretical Network Analysis (GRETNA) toolbox [^3] https://www.nitrc.org/projects/gretna/ and SPM12444https://www.fil.ion.ucl.ac.uk/spm/software/spm12/. The brain is then parcellated using the Automated Anatomical Labeling (AAL) atlas, which defines $116$ ROIs as nodes. For each ROI, we extract the BOLD time series and compute pairwise Pearson correlations [22] between ROI signals to weight their functional connectivity, resulting in a weighted functional brain network for each subject.
@@
We compare SAM-BG with several representative baselines, including supervised and SSL methods. For supervised learning, we adopt BrainGNN[12] and BrainGB[4], where BrainGNN employs an ROI-aware pooling mechanism, and BrainGB provides a standardized benchmark for brain graph analysis. For SSL, we include four competitive methods: GraphCL[28], GCA[33], MA-GCL[8], and CCA-SSG[31]. GraphCL explores several graph augmentation strategies; GCA improves this by adapting to salient graph structure; MA-GCL generates views based on learned embeddings; and CCA-SSG uses CCA to capture perturbation-invariant features without contrastive pairs. For a fair comparison, we adopt the official settings for BrainGNN and BrainGB. All SSL baselines use their original implementations with edge perturbation and 20% labeled data for fine-tuning, simulating low-supervision scenarios.
@@
Our model is implemented using PyTorch and trained on an NVIDIA Tesla P100 GPU. During pre-training, we leverage 20% of the labeled data to guide masker training. For downstream tasks, the dataset is split into 80% for fine-tuning, 10% for validation, and 10% for testing, with the test set strictly excluded from any supervised training during pre-training. The hyperparameters are configured as follows: the MI trade-off coefficient $\beta$ is set to 0.01 for ABIDE and 0.001 for ADHD, while the weight of the decorrelation loss term $\lambda$ is set to $1\text{e}^{-4}$ for ABIDE and 0.001 for ADHD. To balance sparsity and structural completeness, the number of submatrices is set to $K=N/2$. The GNN encoder adopts graph isomorphism networks with bilinear second-order pooling to enhance node interaction and reduce parameters. We conduct 5-fold cross-validation and repeat experiments five times with different random seeds. Results are reported as mean and standard deviation of accuracy (ACC), the area under the ROC curve (AUC), recall, and F1-score.
@@ 4.2 Experimental Results and Discussion
Table 1 presents the complete comparative results between our method and the baseline methods on the ABIDE and ADHD datasets. Overall, supervised learning methods tend to outperform most SSL methods in terms of accuracy, highlighting the importance of label information. Among them, BrainGB attains the highest F1-score, indicating strong robustness under full supervision. Despite this, CCA-SSG achieves competitive results across several metrics, benefiting from its non-contrastive objective, which facilitates better preservation of structural-semantic information. In particular, our proposed method, SAM-BG, consistently outperforms both state-of-the-art supervised and SSL methods across the two datasets. On ABIDE, it improves ACC by 4.1%, AUC by 3.4%, and F1-score by 3.7% compared to the best baseline, although its recall is slightly lower than A-GCL. On ADHD, it achieves 3.3% higher ACC, 0.3% higher recall, and 3.7% higher F1-score, with a marginally lower AUC than A-GCL. Notably, SAM-BG demonstrates more substantial improvements on the ABIDE dataset, suggesting that disease-related connectivity patterns may be more discriminative in autism brain networks. We attribute the superiority of SAM-BG to its two-stage training strategy, which uses limited labeled samples to guide structurally meaningful representation learning on large-scale, unlabeled brain networks.
@@
The Effectiveness of Structure-aware Graph Augmentation Strategy in Data-efficient Settings.
@@ 5 Conclusion
We propose SAM-BG, a novel framework that leverages structural priors to enable biologically meaningful brain graph representation learning under limited supervision. SAM-BG first trains an edge masker on a small labeled subset to extract salient connections, capturing key structural semantics. It then generates structure-aware augmentations by perturbing only non-essential edges, preserving semantic integrity for robust SSL. Experimental results demonstrate that SAM-BG substantially improves psychiatric disorder detection in low-data scenarios. Moreover, it provides explainable insights into brain connectivity patterns, underscoring its strong potential for real-world neuroscience applications.